GATE प्रश्न और उत्तर का अभ्यास करें

प्र: Find the total numbers greater than 4000 that can be formed with digits 2, 3, 4, 5, 6 no digit being repeated in any number ? 2048 0

  • 1
    120
    सही
    गलत
  • 2
    256
    सही
    गलत
  • 3
    192
    सही
    गलत
  • 4
    244
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 3. "192"
व्याख्या :

Answer: C) 192 Explanation: We are having with digits 2, 3, 4, 5 & 6 and numbers greater than 4000 are to be formed, no digit is repeated.   The number can be 4 digited but greater than 4000 or 5 digited.   Number of 4 digited numbers greater than 4000 are4 or 5 or 6 can occupy thousand place => 3 x P34 = 3 x 24 = 72.    5 digited numbers = P55 = 5! = 120  So the total numbers greater than 4000 = 72 + 120 = 192

  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 4. "a, b & c follows"
व्याख्या :

Answer: D) a, b & c follows Explanation: Keeping in mind the safety and convenience of passengers, both 'a' and 'b' follow. 'c' clearly suggests a remedy to the problem and hence it also follows.

प्र: One lady has 2 children, one of her child is boy, what is the probability of having both are boys ? 2046 0

  • 1
    1/3
    सही
    गलत
  • 2
    1/2
    सही
    गलत
  • 3
    2/3
    सही
    गलत
  • 4
    2/5
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 1. "1/3"
व्याख्या :

Answer: A) 1/3 Explanation: In a family with 2 children there are four possibilities:1) the first child is a boy and the second child is a boy (bb)2) the first child is a boy and the second child is a girl (bg)3) the first child is a girl and the second child is a boy (gb)4) the first child is a girl and the second child is a girl (gg) But already given that one child is boy. So we have three possibilities of (bb)(bg)(gb).n(E)= both are boys=BB=1n(S)= 3Required probability P = n(E)/n(S) = 1/3.

प्र: Where are Maglev trains used? 2044 0

  • 1
    London
    सही
    गलत
  • 2
    Russia
    सही
    गलत
  • 3
    Germany
    सही
    गलत
  • 4
    Japan
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 4. "Japan"
व्याख्या :

Answer: D) Japan Explanation: Maglev trains are used in Japan, South Korea and China only. Maglev derived from magnetic levitation using the principle of magnetic repulsion is a system of train transportation that uses two sets of magnets, one set to repel and push the train up off the track as in levitation then another set to move the 'floating train' ahead at great speed taking advantage of no friction. Maglev trains are therefore quieter and smoother than conventional trains, and have the potential for much higher speeds Despite over a century of research and development, maglev transport systems are in operation in just three countries Japan, South Korea and China.   Hence Maglev -- Magnetic-levitation.

प्र: In 1990 a person is 15 years old. In 2000 that same person is 5 years old. How can this be possible? 2044 0

  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर :
व्याख्या :

It can be possible in only B.C. i.e, Before Christ. That was about 2000 years ago. Given In 1990 B.C = 2000 + 1990 = 3990 years ago, a person is 15 years Old => In 1995 B.C = 2000 + 1995 = 3995 years ago, the same person will be 10 years Old Therefore, In 2000 B.C = 2000 + 2000 = 4000 years ago, the same person will be 5 years Old.

प्र: Give the iupac name for the following compound? 2043 0

  • 1
    Sec-Butylcyclohexine
    सही
    गलत
  • 2
    Isopropylcyclohexane
    सही
    गलत
  • 3
    2-Cyclohexybutane
    सही
    गलत
  • 4
    None of the above
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 2. "Isopropylcyclohexane"
व्याख्या :

Answer: B) Isopropylcyclohexane Explanation:

प्र: In a box, there are four marbles of white color and five marbles of black color. Two marbles are chosen randomly. What is the probability that both are of the same color? 2042 0

  • 1
    2/9
    सही
    गलत
  • 2
    5/9
    सही
    गलत
  • 3
    4/9
    सही
    गलत
  • 4
    0
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 3. "4/9"
व्याख्या :

Answer: C) 4/9 Explanation: Number of white marbles = 4 Number of Black marbles = 5 Total number of marbles = 9 Number of ways, two marbles picked randomly = 9C2 Now, the required probability of picked marbles are to be of same color = 4C2/9C2  +  5C2/9C2 = 1/6 + 5/18 = 4/9.

प्र: In a mixture of 240 lt. water is 20% and rest is Milk. What quantity of mixture should be taken out and replaced with water so that water becomes 40%? 2041 0

  • 1
    60 lit
    सही
    गलत
  • 2
    55 lit
    सही
    गलत
  • 3
    45 lit
    सही
    गलत
  • 4
    50 lit
    सही
    गलत
  • उत्तर देखेंउत्तर छिपाएं
  • Workspace

उत्तर : 1. "60 lit"
व्याख्या :

Answer: A) 60 lit Explanation:

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई

      त्रुटि की रिपोर्ट करें

    कृपया संदेश दर्ज करें
    त्रुटि रिपोर्ट सफलतापूर्वक जमा हुई